Can you please help me solve using the elimination method?
8x - 5y = 8
2x + 3y = 2

Answer by 303795(562)

Equation 1 : 8x - 5y = 8
Equation 2 : 2x + 3y = 2
Multiply the second equation by 4 to get
Equation 2 : 8x + 12y = 8
Subtract the original equation 1 from the new equation 2 to get 17y = 0.
y = 0 so substitute that value in one of the original equations
8x - 5y = 8
8x-5*0 =8
8x = 8
so x=1
x = 1 and y = 0